# Fibered ribbon pretzels

### v1

@@ 3. Pretzel knots, graphs and Donaldson’s obstruction
Given a pretzel link $L=P(p_{1},...,p_{n})$, we start this section describing how to associate to $L$ a star-shaped graph $\Gamma_{L}$ which encodes the plumbing instructions to build a 4-manifold $X_{\Gamma_{L}}$, whose boundary $Y_{L}$ is the double cover of $S^{3}$ branched over $L$ (details of the plumbing construction can be found for example in [GS23]). In what follows, when $L$ is clear, we will drop it from the notation. For a given $L$, there are many graphs $\Gamma$ with the described properties. Among them, at most one will encode a negative definite $X_{\Gamma}$ (and precisely one if $L$ is a knot. This is the standard form defined in [NR06, Theorem 5.2]). As explained further below, this negative definite manifold will play a crucial role in our study of sliceness of pretzel knots.
The first step is to associate to $L=P([1^{\pm d}],p_{d+1},\dots,p_{n})$ the plumbing graph in Figure 2, with a central vertex of weight $\mp d$ connected to $n-d$ vertices, each of weight $p_{i}$.
Figure 2. Plumbing graph associated to $L=P([1^{\pm d}],p_{d+1},\dots,p_{n})$.
@@
is negative [NR06, Theorem 5.2]. If $e(Y)$ is positive, then $e(-Y)$ will be negative. The manifold $-Y$ is the double cover of $S^{3}$ branched over the mirror of $L$, which is the pretzel knot with parameters opposite to those of $L$. Conveniently, the properties we are interested in in this article, fiberedness and sliceness/ribbonness of knots, are invariant under mirror imaging, so, up to considering the mirror of $L$, we can always find a negative 4-manifold $X$ to work with. This negative definite manifold will be constructed as follows:
@@
    If $e(Y)$ is negative, then we obtain a graph $\Gamma$ defining a negative definite $X_{\Gamma}$ from the one in Figure 2 by substituting every vertex with positive weight $p_{i}>2$ with a linear graph of $p_{i}-1$ vertices of weight $-2$ connected to the central vertex. For every such substitution, we subtract $1$ from the weight of the central vertex.
@@
    If $e(Y)$ is positive, we change all the signs of the weights in the graph in Figure 2 and then proceed as in the previous point.
This procedure is just an algorithmic description of a series of blow-ups and blow-downs which do not change $Y_{L}$ and yield a negative definite plumbed 4-manifold $X_{\Gamma}$ (see [NR06] for a proof). An example is worked out in Figure 3.
Figure 3. Example of construction of canonical negative definite graph associated to $P(-1,-1,2,3,.-5)$. From left to right, we follow the steps for $e(Y_{L})$ positive.
@@
The reason for constructing a negative definite 4-manifold bounded by the double cover of $S^{3}$ branched over a pretzel knot is that we want to use Donaldson’s Theorem A [Don87] to obstruct sliceness. This obstruction has been described many times in the literature and, in the particular form we will be using it, it was introduced in [Lis07a]. If a knot $K$ is slice (or a link is $\chi$-slice), then its double branched cover $Y_{K}$ bounds a rational homology ball $W$. If $Y_{K}$ also bounds a negative definite manifold $X$, we can construct a closed negative definite 4-manifold as $C:=X\cup_{Y_{K}}-W$. Donaldson’s Theorem A tells us then that the intersection form $Q_{C}$ is diagonalizable. In other words, the lattice $(H_{2}(C),Q_{C})$ is lattice isomorphic to $(\mathbb{Z}^{b_{2}(C)},-\mathrm{Id})$. Moreover, since $W$ is a rational homology ball, a straightforward application of the Mayer-Vietoris sequence with rational coefficients shows that $H_{2}(X)$ includes into $X_{2}(C)$ and, in particular $b_{2}(X)=b_{2}(C)$. This in turn implies that there is a lattice embedding of $(\mathbb{Z}^{b_{2}(X)},Q_{X})$ into $(\mathbb{Z}^{b_{2}(C)},-\mathrm{Id})$, that is, there is an injection $\iota:\mathbb{Z}^{b_{2}(X)}\rightarrow\mathbb{Z}^{b_{2}(C)}$ such that $Q_{X}(a,b)=-\mathrm{Id}(\iota(a),\iota(b))$. The existence of this injection is derived from the existence of a slice disk for $K$. In the proofs in the next section we will show that certain families of pretzel knots are not slice by obstructing the existence of $\iota$.
@@
  • Some times we will look at a sublattice spanned by a subset $S$ of the $e_{i}$ and consider the associated graph (which, importantly, might not be a subgraph of $\Gamma$). This graph with an embedding is obtained by considering only the vertices of $\Gamma$ where the basis vectors in $S$ appear. These vertices will have a new embedding, obtained by simply deleting from their original embedding the vectors not in $S$. The edges of the new graph will be such that they reflect the pairings of the basis vectors in $S$. A simple example is given in Figure 4.
    Figure 4. Example of an embedding of a graph into $\mathbb{Z}^{5}$ and the associated graph with embedding spanned by the subset $S=\{e_{2},e_{3}\}$.
@@
There is an easy algorithm to determine the Wu set, that is the set of vertices $v_{i}^{w}$ of the graph which satisfy $w=\sum_{i}v_{i}^{w}$ [see for example [Lec15, Section 3.2] and [NR06]]. An important feature that we want to note here is that, for ribbon pretzel knots and when working with a negative definite graph, we can always assume that the embedding of the Wu class in the closed manifold $X_{\Gamma}\cup W$ (where $W$ is a rational homology ball) is the sum of all the basis vectors of the standard negative definite lattice, in symbols, $\iota(w)=\sum_{i}^{k}e_{i}$. Indeed, in this case, Equation (2) via the embedding $\iota$ reads
$-\mathrm{Id}(\iota(w),x)\equiv-\mathrm{Id}(x,x)\mod 2\quad\quad\forall x\in\mathrm{H_{2}}(X_{\Gamma}\cup W;\mathbb{Z}).$
Considering this last congruence in particular for the classes $x=e_{i}$, for each standard basis vector $e_{i}$, we obtain that all the coefficients appearing in the expression of $\iota(w)$ need to be odd (and in particular non-zero). Moreover, since ribbon knots have signature zero, Equation (1) reads $k=Q_{\Gamma}(w,w)=-\mathrm{Id}(\iota(w),\iota(w))$, which immediately implies that the coefficients need to be at most 1 in absolute value. By eventually applying some automorphisms of $(\mathbb{Z}^{k},-\mathrm{Id})$, sending a basis element $e_{i}$ to $-e_{i}$, we can assume all the coefficients are $+1$ and therefore $\iota(w)=\sum_{i}^{k}e_{i}$ as claimed.
@@ 4. Proofs
The pretzel knot ±$\pm P(1,1,1,1,-3,-3,-3)=\pm 10_{75}$ is the only Type 1 fibered ribbon pretzel knot.
@@
From Gabai’s classification (Theorem 2.1) of fibered pretzel knots, we know that Type 1 fibered pretzel knots are of the form $K=P(1,\dots,1,-3,\dots,-3)$ or their mirrors $\bar{K}=P(-1,\dots,-1,3,\dots,3)$ with at least one unitary parameter and an odd number of parameters. Our task is to determine which among these are ribbon and it suffices to study the family of knots $K$. To this end, we will associate to $K$ a negative definite graph and we will try to embed it into the standard negative lattice of the same rank (see Sections 2 and 3 for notation and strategy).
The negative definite graphs associated to $K$, together with the Wu classes, are depicted in Figure 5.
Figure 5. Negative definite graphs associated to $K$. The Wu-class in each case is depicted in red.
@@
We assume $K$ is slice. Then $\sigma(K)=0$ and equation (1) reads $0=-(n-d+1)+d$ for the top graph and $0=-(2(n-d)+1)-2d+2d-n$ for the bottom graph in Figure 5. The latter condition immediately yields a contradiction: $3n+1=2d$ is incompatible with $n>4d$, the negative definite condition of the graph. As such, we will restrict attention to graphs of the first kind with the constraint $n+1=2d$.
@@
Since $K$ is slice, the lattice $(\mathbb{Z}^{\mathrm{rk}(\Gamma)},Q_{\Gamma})$ embeds into the standard negative lattice of the same rank. With the notation conventions of Section 3, we will show that this only happens for the knot $10_{75}$. We will argue depending on the value of $m:=n-d$, which in our case can be written $m=d-1$.
@@
Notice that for $m\in\{1,2\}$ the manifold $Y_{K}$ is a lens space. In these two cases, while it is immediate to show by hand that there is no embedding, the non-existence follows from work of Lisca [Lis07a].
@@
When $m=3$, our constraint implies $d=4$, corresponding to the knot $10_{75}$ and an embedding into the diagonal negative definite lattice of rank 4 is as follows. The central vertex $v_{c}$, which is the Wu class, has embedding $v_{c}=e_{1}+e_{2}+e_{3}+e_{4}$. The three weight 3 vertices have embeddings $v_{1}=-e_{1}+e_{4}-e_{2}$, $v_{2}=-e_{1}+e_{3}-e_{4}$ and $v_{3}=-e_{1}+e_{2}-e_{3}$. The manifold $Y_{K}$ and this embedding are well known and studied (see for example [SSW08, Figure 23]).
@@
If $m>4$, we claim there is no embedding. We are trying to embed $m$ weight $-3$ orthogonal vertices and the central vertex into $(\mathbb{Z}^{m+1},-\mathrm{Id})$. We will argue that the graphs we are considering embed only in standard lattices of rank strictly larger than $m+1$. Up to sign, the only embedding of a weight -3 vertex is of the form $v=e_{\alpha}+e_{\beta}+e_{\gamma}$. It is immediate to see that the minimum number of basis vectors needed to embed:
  • one weight -3 vertex is three, $v=e_{\alpha}+e_{\beta}+e_{\gamma}$;
  • two orthogonal weight -3 vertices is four, $v_{1}=e_{\alpha}-e_{\beta}+e_{\gamma}$ and $v_{2}=e_{\alpha}+e_{\beta}+e_{\delta}$;
  • three orthogonal weight -3 vertices is four, $v_{1}=e_{\alpha}-e_{\beta}+e_{\gamma}$, $v_{2}=e_{\alpha}+e_{\beta}+e_{\delta}$ and $v_{3}=e_{\alpha}-e_{\gamma}-e_{\delta}$.
@@
Notice that, once the embedding of a set of three orthogonal weight -3 vertices is fixed, any further weight -3 orthogonal vertex cannot share any basis vectors with these three. (Any two such vertices must share either zero or two basis vectors in their embedding since they are orthogonal.) It follows easily that when embedding $n$ weight -3 orthogonal vertices into a rank $k$ lattice, the difference $k-n$ is minimal when $n=1\mod 3$, in which case $k-n=n/3$, so in particular is greater than 1 whenever $n>3$.
@@
To finish the proof, we need to show that the knot $10_{75}$ is indeed ribbon. While this information is available in KnotInfo [LMar], for completeness and comparison, we include the band moves that turn this knot into a disjoint union of two unknots, both in the minimal alternating projection of $10_{75}$ and in the standard projection of this knot in pretzel form. See Figure 6. ∎
@@
Figure 6. To the left, a pretzel projection of $K=P([1^{4}],-3,-3,-3)$ with a band describing a move to two unlinked unknots. To the right, the same band after isotopying $K$ to an alternating projection that allows us to identify it as the knot $10_{75}$.
@@
We move on to analyze pretzel knots of Types 2 and 3. From the point of view of sliceness/ribonness these were thoroughly studied in [Lec15], but in that article the parameters ±$\pm 1$ were excluded from the discussion. We divide the rest of this section in two. The first subsection will deal with the case of no ±$\pm 1$ parameters and collect known facts about this class. The second subsection will deal with the remaining cases which need to be studied from scratch from the sliceness perspective.
@@
We are not aware of anywhere in the literature where the results from Gabai’s classification of pretzel knots of Types 2 and 3, Theorem 2.1, have been contrasted with [Lec15, Theorem 1.1]. We state here what we obtain when comparing these two classification results.
@@
    Type 2A fibered pretzel knots are never slice. Indeed, all Type 2A pretzel knots are of the form (p2) or (p3) in [Lec15]. In that article, Lemmas 5.4 and 5.5 establish, under the only assumption of these knots being slice, that the number of positive and negative parameters must differ exactly by one, whereas Type 2A requires this difference to be two.
@@
    Type 3A fibered pretzel knots cannot be slice. Indeed, these pretzels have an unequal number of positive and negative parameters which excludes them from being slice by [Lec15, Lemma 5.3]. (Note that Type 3 pretzel knots are of the form (p1).)
    Type 3B pretzel knots with no unitary parameters are never fibered: since the number of positive and negative parameters is equal, $L^{\prime}$ has an even number of entries; however, a Type 3B fibered pretzel knot has an asscociated $L^{\prime}$ with an odd number of entries.
    Finally, Type 3C pretzel knots are fibered if and only if there is a unique parameter of minimal absolute value. From the ribbon point of view and up to reordering, Type 3 ribbon pretzel knots are determined by a set of parameters which consists of a certain number of paired up parameters of the same absolute value and opposite sign, and a last unordered pair of the form ±$\pm(k,-k-1)$ with $k>1$. It follows that in order to be Type 3 fibered and ribbon, the parameter $k$ needs to be the unique minimal positive term. For example $P(-3,3,-3,4)$ is ribbon but not fibered, while $P(-3,3,-3,2)$ is ribbon and fibered.
@@
    $K=\pm P(k,-k-1,q_{1},\ldots,q_{r},-q_{1}\ldots,-q_{r}$) with $q_{i}\geq 3$, $r\geq 0$, and $1<k<q_{i}$ for all $i$.
Some further immediate remarks regarding the order of the parameters are pertinent. When the order is relevant to fiberedness, the condition is generally that we need parameters alternating sign. When we are dealing with ribonness, we need to be able to pair up the parameters of opposite sign and equal absolute value. For example, the Type 2B detectably ribbon knot $K=P(5,3,7,-7,-3,-5,2)$ is not fibered, while its concordant mutant $K^{\prime}=P(5,-3,7,-7,3,-5,2)$ is fibered (and ribbon). Finally, the mutant $K^{\prime\prime}=P(7,-3,5,-7,3,-5,2)$, which is not detectably ribbon (and which we conjecture is not slice) is fibered, while the mutant $K^{\prime\prime\prime}=P(7,3,5,-7,-3,-5,2)$ is niether fibered not ribbon (conjecturally).
@@
The existence of an embedding $\iota$ of $(\Gamma_{1},Q_{\Gamma_{1}})$ into the standard negative lattice of the same rank and the particular form of the embedding of the Wu class, imply that the central vertex of $\Gamma_{1}$, call it $v_{c}$, has weight at most $-r-1$ (each of the the $r+1$ vertices in the Wu set shares at least one basis vector with the central vertex, and these are all different). On the other hand, since $K$ is fibered, $d+t-r=\pm 2$ and therefore, the weight of the central vertex satisfies $v_{c}\cdot v_{c}=-d-t-1=-r-1\mp 2$. As this quantity is bounded above by $-r-1$, the sign of 2 is fixed, giving $v_{c}\cdot v_{c}=-r-3$, and $d+t-r=2$. Moreover, Equation (1) in this context reads:
@@
Labelling the vertex with weight $n_{i}$ by $v_{i}$ , and $w^{\prime}$ the unique vertex in the Wu set with weight $-2$, we can fix the embedding of the Wu class as the sum of $v_{i}=e^{i}_{1}+\dots+e^{i}_{|n_{i}|}$ and $w^{\prime}=f_{1}+f_{2}$, where the standard basis of the negative diagonal lattice is given by $\{e^{1}_{1},\dots,e^{r}_{|n_{r}|},f_{1},f_{2}\}$ and, via Equation (3), the rank of $\Gamma_{1}$ is $\sum|n_{i}|+2$. We can moreover fix the embedding $v_{c}=-e^{1}_{1}-e^{2}_{1}-\dots-e^{r}_{1}-f_{1}+e^{1}_{2}-e^{1}_{3}$: we need the first $r+1$ basis vectors with these signs to connect the central vertex to the $v_{i}$ and to $w^{\prime}$; the last two basis vectors yield the correct weight for $v_{c}$. Since $|n_{i}|\geq 3$, up to relabelling the vertices, we can pick the two basis vectors from the embedding of $v_{1}$. The only real choice we have made in the embedding of $v_{c}$ is the sign of $e^{1}_{2}$ (which forces the sign of $e^{1}_{3}$). Finally, we label the $-2$-weight vertices arising from blowing up/down the $p_{j}$-weight vertex of $\Gamma_{K}$ by $u^{j}_{1},\ldots,u^{j}_{p_{j}-1}$, where sub-indices increase moving away from $v_{c}$, denote by $U^{j}$ the union of these vertices, which we will refer to as a $-2$-chain, and define $U=\cup_{j}U^{j}$.

### v2

@@ 3. Pretzel knots, graphs and Donaldson’s obstruction
Given a pretzel link $L=P(p_{1},...,p_{n})$, we start this section describing how to associate to $L$ a star-shaped graph $\Gamma_{L}$ which encodes the plumbing instructions to build a 4-manifold $X_{\Gamma_{L}}$, whose boundary $Y_{L}$ is the double cover of $S^{3}$ branched over $L$ (details of the plumbing construction can be found for example in [gompf20234]). In what follows, when $L$ is clear, we will drop it from the notation. For a given $L$, there are many graphs $\Gamma$ with the described properties. Among them, at most one will encode a negative definite $X_{\Gamma}$ (and precisely one if $L$ is a knot. This is the standard form defined in [NeuRay, Theorem 5.2]). As explained further below, this negative definite manifold will play a crucial role in our study of sliceness of pretzel knots.
The first step is to associate to $L=P([1^{\pm d}],p_{d+1},\dots,p_{n})$ the following plumbing graph, with a central vertex of weight $\mp d$ connected to ndn-d vertices, each of weight $p_{i}$.
[…]
@@
is negative [NeuRay, Theorem 5.2]. If $e(Y)$ is positive, then e(Y)e(-Y) will be negative. The manifold Y-Y is the double cover of $S^{3}$ branched over the mirror of $L$, which is the pretzel knot with parameters opposite to those of $L$. Conveniently, the properties we are interested in in this article, fiberedness and sliceness/ribbonness of knots, are invariant under mirror imaging, so, up to considering the mirror of $L$, we can always find a negative 4-manifold $X$ to work with. This negative definite manifold will be constructed as follows:
@@
    If $e(Y)$ is negative, then we obtain a graph $\Gamma$ defining a negative definite $X_{\Gamma}$ from the one above by substituting every vertex with positive weight $p_{i}>2$ with a linear graph of $p_{i}-1$ vertices of weight 2-2 connected to the central vertex. For every such substitution, we subtract $1$ from the weight of the central vertex.
@@
    If $e(Y)$ is positive, we change all the signs of the weights in the graph above and then proceed as in the previous point.
This procedure is just an algorithmic description of a series of blow-ups and blow-downs which do not change $Y_{L}$ and yield a negative definite plumbed 4-manifold $X_{\Gamma}$ (see [NeuRay] for a proof). For example, starting with the pretzel knot $L=P(-1,-1,2,3,-5)=P([1^{-2}],2,3,-5)$, we compute the associated Euler number
$e(Y_{L})=2-\frac{1}{2}-\frac{1}{3}+\frac{1}{5}>0,$
and then follow the steps described in (2) above which are illustrated with the following graphs, yielding the rightmost one, which is the one producing a negative definite 4-manifold via plumbing.
[…]
@@
The reason for constructing a negative definite four manifold bounded by the double cover of $S^{3}$ branched over a pretzel knot is that we want to use Donaldson’s Theorem A [Donaldson] to obstruct sliceness. This obstruction has been described many times in the literature and, in the particular form we will be using it, it was introduced in [lisca]. If a knot $K$ is slice (or a link is $\chi$-slice), then its double branched cover $Y_{K}$ bounds a rational homology ball $W$. If $Y_{K}$ also bounds a negative definite manifold $X$, we can construct a closed negative definite 4-manifold as $C:=X\cup_{Y_{K}}-W$. Donaldson’s Theorem A tells us then that the intersection form $Q_{C}$ is diagonalizable. In other words, the lattice $(H_{2}(C),Q_{C})$ is lattice isomorphic to $(\mathbb{Z}^{b_{2}(C)},-\mathrm{Id})$. Moreover, since $W$ is a rational homology ball, a straightforward application of the Mayer-Vietoris sequence with rational coefficients shows that $H_{2}(X)$ injects into $X_{2}(C)$ and, in particular $b_{2}(X)=b_{2}(C)$. This in turn implies that there is a lattice embedding of $(\mathbb{Z}^{b_{2}(X)},Q_{X})$ into $(\mathbb{Z}^{b_{2}(C)},-\mathrm{Id})$, that is, there is an injection $\iota:\mathbb{Z}^{b_{2}(X)}\rightarrow\mathbb{Z}^{b_{2}(C)}$ such that $Q_{X}(a,b)=-\mathrm{Id}(\iota(a),\iota(b))$. The existence of this injection is derived from the existence of a slice disk for $K$. In the proofs in the next section we will show that certain families of pretzel knots are not slice by obstructing the existence of $\iota$.
@@
  • Some times we will look at a sublattice spanned by a subset $S$ of the $e_{i}$ and consider the associated graph (which, importantly, might not be a subgraph of $\Gamma$). This graph with an embedding is obtained by considering only the vertices of $\Gamma$ where the basis vectors in $S$ appear. These vertices will have a new embedding, obtained by simply deleting from their original embedding the vectors not in $S$. The edges of the new graph will be such that they reflect the pairings of the basis vectors in $S$. A simple example illustrates the idea: in the following diagram, to the left we have an embedding of a graph into the lattice ($\mathbb{Z}^{5}=\langle e_{1},\dots,e_{5}\rangle,-\mathrm{Id})$; letting $S$ be the subset $\{e_{2},e_{3}\}$ we can consider the associated graph, which is given to the right of the diagram.
[…]
@@
There is an easy algorithm to determine the Wu set, that is the set of vertices $v_{i}^{w}$ of the graph which satisfy $w=\sum_{i}v_{i}^{w}$ [see for example [pretzel, Section 3.2] and [NeuRay]]. An important feature that we want to note here is that, for ribbon pretzel knots and when working with a negative definite graph, we can always assume that the embedding of the Wu class in the closed manifold $X_{\Gamma}\cup W$ (where $W$ is a rational homology ball) is the sum of all the basis vectors of the standard negative definite lattice, in symbols, $\iota(w)=\sum_{i}^{k}e_{i}$. Indeed, as $\det(Y_{K})=\det(Q_{\Gamma})$ is odd, $\iota(H_{2}(X_{\Gamma}))\subset H_{2}(X_{\Gamma}\cup W)$ has odd index, so reduction modulo 2 is surjective; thus in this case, Equation (2) extends to all $x\in H_{2}(X_{\Gamma}\cup W)$ and via the embedding $\iota$ reads
$\mathrm{Id}(\iota(w),x)\equiv-\mathrm{Id}(x,x)\mod 2\quad\quad\forall x\in\mathrm{H_{2}}(X_{\Gamma}\cup W;\mathbb{Z}).$
Considering this last congruence in particular for the classes $x=e_{i}$, for each standard basis vector $e_{i}$, we obtain that all the coefficients appearing in the expression of $\iota(w)$ need to be odd (and in particular non-zero). Moreover, since ribbon knots have signature zero, Equation (1) reads $k=Q_{\Gamma}(w,w)=-\mathrm{Id}(\iota(w),\iota(w))$, which immediately implies that the coefficients need to be at most 1 in absolute value. By eventually applying some automorphisms of $(\mathbb{Z}^{k},-\mathrm{Id})$, sending a basis element $e_{i}$ to $e_{i}$, we can assume all the coefficients are +1+1 and therefore $\iota(w)=\sum_{i}^{k}e_{i}$ as claimed.
@@ 4. Proofs
The pretzel knot $\pm P(1,1,1,1,-3,-3,-3)=\pm 10_{75}$ is the only Type 1 fibered ribbon pretzel knot.
@@
From Gabai’s classification (Theorem 2.1) of fibered pretzel knots, we know that Type 1 fibered pretzel knots are of the form $K=P(1,\dots,1,-3,\dots,-3)$ or $\bar{K}=P(-1,\dots,-1,3,\dots,3)$ with at least one unitary parameter and an odd number of parameters. Our task is to determine which among these are ribbon and it suffices to study the family of knots $K$. To this end, we will associate to $K$ a negative definite graph and we will try to embed it into the standard negative lattice of the same rank (see Sections 2 and 3 for notation and strategy).
In order to determine which is the negative definite graph associated to a Type 1 fibered knot with $n$ parameters $K=([1^{d}],-3,\dots,-3)$ we need to first compute and understand the sign of its Euler number
$e(Y_{K})=\frac{n-d}{3}-d.$
If $n<4d$, then $e(Y_{K})<0$ and the negative definite graph $\Gamma_{K}$, with the Wu-set marked in red, is
$\bullet3-3$$d\bullet\bullet\bullet3-3\bullet$
On the other hand, if $n>4d$, then $e(Y_{K})>0$ and the negative definite graph $\Gamma^{\prime}_{K}$ with its corresponding Wu-set in red is
$d-n\bullet\bullet\bullet\bullet\bullet\bullet\bullet\bullet\bullet2-22-22-22-2$2-2
@@
We assume $K$ is slice. Then $\sigma(K)=0$ and Equation (1) reads 0=(nd+1)+d0=-(n-d+1)+d for $\Gamma_{K}$ and 0=(2(nd)+1)+2(nd)2d+n0=-(2(n-d)+1)+2(n-d)-2d+n for $\Gamma^{\prime}_{K}$. The latter condition immediately yields a contradiction: $n=2d+1$ is incompatible with $n>4d$, the negative definite condition of the graph. As such, we will restrict attention to graphs of the first kind with the constraint $n+1=2d$.
@@
Since $K$ is slice, the lattice $(\mathbb{Z}^{\mathrm{rk}(\Gamma)},Q_{\Gamma})$ embeds into the standard negative lattice of the same rank. With the notation conventions of Section 3, we will show that this only happens for the knot $10_{75}$. We will argue depending on the value of m:=ndm:=n-d, which in our case can be written m=d1m=d-1.
@@
Notice that for $m\in\{1,2\}$ the manifold $Y_{K}$ is a lens space. In these two cases, while it is immediate to show by hand that there is no embedding, the non-existence follows from work of Lisca [lisca].
@@
When $m=3$, our constraint implies $d=4$, corresponding to the knot $10_{75}$ and an embedding into the diagonal negative definite lattice of rank 4 is as follows. The central vertex $v_{c}$, which is the Wu class, has embedding $v_{c}=e_{1}+e_{2}+e_{3}+e_{4}$. The three weight 3 vertices have embeddings $v_{1}=-e_{1}+e_{4}-e_{2}$, $v_{2}=-e_{1}+e_{3}-e_{4}$ and $v_{3}=-e_{1}+e_{2}-e_{3}$. The manifold $Y_{K}$ and this embedding are well known and studied (see for example [Stipsicz, Figure 23]).
@@
If $m>3$, we claim there is no embedding. We are trying to embed $m$ weight 3-3 orthogonal vertices and the central vertex into $(\mathbb{Z}^{m+1},-\mathrm{Id})$. We will argue that the graphs we are considering embed only in standard lattices of rank strictly larger than $m+1$. Having embedded the central vertex $v_{c}=e_{1}+e_{2}+\cdots+e_{m+1}$, the only embedding of a linked weight 3-3 vertex is of the form $v=-e_{i}+e_{j}-e_{k}$. It is immediate to see that the minimum number of basis vectors needed to embed:
  • one weight 3-3 vertex is three, $v_{1}=-e_{1}+e_{2}-e_{3}$;
  • two orthogonal weight 3-3 vertices is four, $v_{1}=-e_{1}+e_{2}-e_{3}$ and $v_{2}=-e_{1}-e_{2}+e_{4}$;
  • three orthogonal weight 3-3 vertices is four, $v_{1}=-e_{1}+e_{2}-e_{3}$, $v_{2}=-e_{1}-e_{2}+e_{4}$, and $v_{3}=-e_{1}+e_{3}-e_{4}$.
@@
Notice that, once the embedding of a set of three orthogonal weight 3-3 vertices using four basis vectors is fixed, any further weight 3-3 orthogonal vertex linked to the central vertex cannot share any basis vectors with these three. Similarly, once a set of $3l$ orthogonal weight 3-3 vertices using $4l$ basis vectors is fixed, any further weight 3-3 orthogonal vertex cannot share any basis vectors with these. It follows easily that when embedding $n$ weight 3-3 orthogonal vertices into a rank $k$ lattice, the difference knk-n is minimal when $n=0\mod 3$, in which case kn=n/3k-n=n/3, so in particular is greater than 1 whenever $n>3$.
@@
To finish the proof, we need to show that the knot $10_{75}$ is indeed ribbon. While this information is available in KnotInfo [knotinfo], for completeness and comparison we include the band moves that turn this knot into a disjoint union of two unknots, both in the minimal alternating projection of $10_{75}$ and in the standard projection of this knot in pretzel form. See Figure 2. ∎
@@
Figure 2. To the left, a pretzel projection of $K=P([1^{4}],-3,-3,-3)$ with a band describing a move to two unlinked unknots. To the right, the same band after isotoping $K$ to an alternating projection that allows us to identify it as the knot $10_{75}$.
@@
We move on to analyze pretzel knots of Types 2 and 3. From the point of view of sliceness/ribonness these were thoroughly studied in [pretzel], but in that article the parameters $\pm 1$ were excluded from the discussion. We divide the rest of this section in two. The first subsection will deal with the case of no $\pm 1$ parameters and collect known facts about this class. The second subsection will deal with the remaining cases which need to be studied from scratch from the sliceness perspective.
@@
We are not aware of anywhere in the literature where the results from Gabai’s classification of pretzel knots of Types 2 and 3, Theorem 2.1, have been contrasted with [pretzel, Theorem 1.1]. We state here what we obtain when comparing these two classification results.
@@
    Type 2A fibered pretzel knots are never slice. Indeed, all Type 2A pretzel knots are of the form (p2) or (p3) in [pretzel]. In that article, Lemmas 5.4 and 5.5 establish, under the only assumption of these knots being slice, that the number of positive and negative parameters must differ exactly by one, whereas Type 2A requires this difference to be two.
@@
    Type 3A fibered pretzel knots cannot be slice. Indeed, these knots have an unequal number of positive and negative parameters, which excludes them from being slice by [pretzel, Lemma 5.3]. (Note that Type 3 pretzel knots are of the form (p1).)
    Type 3B pretzel knots with no unitary parameters are never fibered: since the number of positive and negative parameters is equal, $L^{\prime}$ has an even number of entries; however, a Type 3B fibered pretzel knot has an associated $L^{\prime}$ with an odd number of entries.
    Finally, Type 3C pretzel knots are fibered if and only if there is a unique parameter of minimal absolute value. From the ribbon point of view and up to reordering, Type 3 ribbon pretzel knots are determined by a set of parameters which consists of a certain number of paired up parameters of the same absolute value and opposite sign, and a last unordered pair of the form $\pm(k,-k-1)$ with $k>1$. It follows that in order to be Type 3C fibered and ribbon, the parameter $k$ needs to be the unique minimal positive term. For example P(3,3,3,4)P(-3,3,-3,4) is ribbon but not fibered, while P(3,3,3,2)P(-3,3,-3,2) is ribbon and fibered.
@@
    $K=\pm P(k,-k-1,q_{1},\ldots,q_{r},-q_{1},\ldots,-q_{r}$) with $q_{i}\geq 3$, $r\geq 0$, and $1<k<q_{i}$ for all $i$.
A final immediate remark regarding the order of the parameters: When the order is relevant to fiberedness, the condition is generally that we need parameters alternating sign; when we are dealing with ribonness, we need to be able to pair up the parameters of opposite sign and equal absolute value. In the introduction we have showcased this fact with the knot P(3,3,5,5,7,7,4)P(3,-3,5,-5,7,-7,4) and its mutants.
@@
The existence of an embedding $\iota$ of $(\Gamma_{1},Q_{\Gamma_{1}})$ into the standard negative lattice of the same rank, and the particular form of the embedding of the Wu class, imply that the central vertex of $\Gamma_{1}$, call it $v_{c}$, has weight at most r1-r-1 (each of the $r+1$ vertices in the Wu set shares at least one basis vector with the central vertex, and these are all different). On the other hand, since $K$ is fibered, $d+t-r=\pm 2$ and therefore, the weight of the central vertex satisfies $v_{c}\cdot v_{c}=-d-t-1=-r-1\mp 2$. As this quantity is bounded above by r1-r-1, the sign of 2 is fixed, giving $v_{c}\cdot v_{c}=-r-3$, and d+tr=2d+t-r=2. Moreover, Equation (1) in this context reads:
@@
Labelling the vertex with weight $n_{i}$ by $v_{i}$, and $w^{\prime}$ the unique vertex in the Wu set with weight 2-2, we can fix the embedding of the Wu class as the sum of $v_{i}=e^{i}_{1}+\dots+e^{i}_{|n_{i}|}$ and $w^{\prime}=f_{1}+f_{2}$, where the standard basis of the negative diagonal lattice is given by $\{e^{1}_{1},\dots,e^{r}_{|n_{r}|},f_{1},f_{2}\}$ and, via Equation (3), the rank of $\Gamma_{1}$ is $\sum|n_{i}|+2$. We can moreover fix the embedding $v_{c}=-e^{1}_{1}-e^{2}_{1}-\dots-e^{r}_{1}-f_{1}+e^{1}_{2}-e^{1}_{3}$: we need the first $r+1$ basis vectors with these signs to connect the central vertex to the $v_{i}$ and to $w^{\prime}$; the last two basis vectors yield the correct weight for $v_{c}$. Since $|n_{i}|\geq 3$, up to relabelling the vertices, we can pick the two basis vectors from the embedding of $v_{1}$. The only real choice we have made in the embedding of $v_{c}$ is the sign of $e^{1}_{2}$ (which forces the sign of $e^{1}_{3}$). Finally, we label the 2-2-weight vertices arising from blowing up/down the $p_{j}$-weight vertex of $\Gamma_{K}$ by $u^{j}_{1},\ldots,u^{j}_{p_{j}-1}$, where sub-indices increase moving away from $v_{c}$, denote by $U^{j}$ the union of these vertices, which we will refer to as a 2-2-chain, and define $U=\cup_{j}U^{j}$.